OpenStudy (anonymous):

There is a container in the shape of a cylinder with hemispheres on both ends The length the entire container is 149 ft Find the largest possible volume

OpenStudy (anonymous):

I got 2x+y=149 Where y is the length of the cylinder And that is y=149-2x The volume of the entire thing is (4/3)(pi)x^3+(pi)(x^2)(y)

OpenStudy (anonymous):

Substituting 149-2x for y in the volume equation gave me (-2/3)x^3(pi)+149(pi)x^2= Volume Not sure what interval to use to find the max by graphing now

OpenStudy (anonymous):

@jdoe0001 could you help me?

OpenStudy (anonymous):

Here's a drawing to the best of my ability